Given that a = 3i - 5j, determine the angle vector a makes with the positive x-axis b. Given that c = 5i + 3j, d = 3i – 2j. Find the magnitude of (i) c + d (ii) c – d

1
Expert's answer
2022-06-16T09:53:59-0400

a.


"\tan\theta=\dfrac{-5}{3}"

Quadrant IV


"\theta=360\degree-\tan^{-1}\dfrac{5}{3}\approx301\degree"

The acute angle which a vector makes with the positive x-axis is "59\degree" clockwise.


b.

(i)

"\vec{c}+\vec{d}=(5+3)\vec{i}+(3-2)\vec{j}"

"=8\vec{i}+\vec{j}"

"|\vec{c}+\vec{d}|=\sqrt{(8)^2+(1)^2}=\sqrt{82}"

(ii)

"\vec{c}-\vec{d}=(5-3)\vec{i}+(3-(-2))\vec{j}"

"=2\vec{i}+5\vec{j}"

"|\vec{c}-\vec{d}|=\sqrt{(2)^2+(5)^2}=\sqrt{29}"